OpenStudy (anonymous):

how do i get 4e alone? 4e-8 = 32

OpenStudy (amistre64):

you remove the -8

OpenStudy (amistre64):

if you have a hole that is 8 feet deep; how do you get rid of the hole?

OpenStudy (anonymous):

Being serious!

OpenStudy (amistre64):

so am i

OpenStudy (shadowfiend):

Jay, please be polite. amistre is trying to drawn an analogy so you understand the way you solve this equation.

OpenStudy (anonymous):

Sorry Amistre

OpenStudy (anonymous):

subtract 8 from both sides

OpenStudy (amistre64):

'sok, no offense taken :)

OpenStudy (anonymous):

Whoops i meant add 8 on both sides

OpenStudy (amistre64):

add 8 is what I would do .... ;)

OpenStudy (anonymous):

ya but like u gotta understand y cuz this is pretty basic

OpenStudy (anonymous):

to left or right side?

OpenStudy (anonymous):

both sides

OpenStudy (amistre64):

both sides have to stay equal; so what you do to one side HAS to be done to the other

OpenStudy (anonymous):

so inorder to remove it from the left side you would have to add 8 right cus -8+8=0

OpenStudy (anonymous):

So would it be 4e 8 - 8 = 32 + 8 ?

OpenStudy (anonymous):

can you put the equation together for mee lol getting confused

OpenStudy (anonymous):

and then you have to add 8 to the other side since whtvr you do to one side you do to the other

OpenStudy (amistre64):

disregarding typos; yes 4e-8 +8 = 32 + 8

OpenStudy (anonymous):

ya that is it

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

A + 18 = -32 ? Its kinda like the other one but different problem

OpenStudy (anonymous):

what you need to isolate A?

OpenStudy (anonymous):

what do u think shld be done

OpenStudy (anonymous):

?? lol

OpenStudy (anonymous):

I just want to see if I explained it clearly and u understood what i said

OpenStudy (anonymous):

cuz like this is pretty imp to understand

OpenStudy (anonymous):

so would it be A + 18 - 18 = -32 + 18?

OpenStudy (anonymous):

Ya good one You deserved a medal!!! 'D

OpenStudy (anonymous):

K but you gotta like clean up the equation though

OpenStudy (anonymous):

A isnt isolated as of yet

OpenStudy (anonymous):

You still gotta finish it off

OpenStudy (anonymous):

so what do we have to do to get it isolated?

OpenStudy (anonymous):

A = 50

OpenStudy (anonymous):

Well just add and subtract the values here

OpenStudy (anonymous):

remember -32+18

OpenStudy (anonymous):

You were pretty close

OpenStudy (anonymous):

-32+18=-14

OpenStudy (anonymous):

so the answer wld be A=-14

OpenStudy (anonymous):

Did u get that?

OpenStudy (preetha):

Jay, I love the way rid613 (and amistre earlier) were walking you through the problem. I think you will be able to handle it better on your own the next time you see this - say in a test. Its not the answer alone - but how you do it and why you do what you do. Thanks rid613 and amistre.

OpenStudy (anonymous):

Yes :) Can you help me get through a couple more?

OpenStudy (anonymous):

ok try them onyour own and I will tell u if the answers are correct ok?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Okay :) M/9 + 26 = 30 would it be M/9 -26 + 26 = 30+ 26

OpenStudy (anonymous):

yes but then we sill have a problem

OpenStudy (anonymous):

jay are you there

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

M/9=56

OpenStudy (anonymous):

M isnt isolated

OpenStudy (amistre64):

i tend to see this easier in an up and down format: M/9 + 26 = 30 -26 -26 -------------- M/9 + 0 = 4

OpenStudy (anonymous):

ya amistre is correct

OpenStudy (anonymous):

See if you write it out the way amistre did it becomes simpler to see and you dont get mixed up

OpenStudy (amistre64):

i get mixed up, which is why i do that ;)

OpenStudy (anonymous):

ok how do u think we wld be able to get rid of that 9?

OpenStudy (anonymous):

-9?

OpenStudy (amistre64):

better yet, how do we get rid of the "/9"

OpenStudy (anonymous):

ya better put it that way!!!!

OpenStudy (anonymous):

well that wldnt work

OpenStudy (amistre64):

look at it as an opposite operation to perform

OpenStudy (anonymous):

would it be -9

OpenStudy (anonymous):

no i will show you what it is

OpenStudy (amistre64):

the opposite of divideing is mulitiplying; so multiply by 9

OpenStudy (anonymous):

\[m/\]